# Rarefied neutral flow in orificed hollow cathodes

### v1

@@ IV.1.2 Subsonic flow
In some cases where $d_{o}=d_{k}$, the keeper induces a (static) pressure rise that is large enough to prevent the orifice from choking. In this configuration the flow remains subsonic throughout the orifice-keeper region, and the supersonic expansion only occurs in the keeper orifice (as opposed to the main cathode orifice). Because the flow is subsonic in this particular case, the keeper-to-total pressure ratio, $P_{ko}/P_{T}$, is higher than that in the supersonic case. Figures 20 and 21 show a contour of the Mach number and static pressure for the subsonic case, respectively. For high-enough mass flow rates, however, the flow may be supersonic. The exact point at which the flow transitions from subsonic to supersonic for the case where $d_{o}=d_{k}$ is beyond the scope of this article.
Figure 16: Centerline values for $d_{o}/d_{c}=0.1$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 17: Centerline values for $d_{o}/d_{c}=0.3$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 18: Centerline values for $d_{o}/d_{c}=0.5$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
Figure 19: Centerline values for $d_{o}/d_{c}=0.7$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$.
@@ V Conclusion
We have performed a parametric study to quantify the effect of the changes in keeper geometrical parameters on the neutral flow field within hollow cathodes. Because cathode ignition occurs typically between the keeper and orifice plates, we focused our study on the keeper-orifice region. Using the DSMC method, we showed that the neutral cathode flow results in most cases in an underexpanded, supersonic, rarefied jet with Knudsen numbers above 0.1, along with a toroidal subsonic recirculation zone. The recirculation zone is bounded by the underexpanded jet viscous layer, the cathode orifice plate, and the keeper plate. We have shown that simple conservation laws can explain the experimentally and numerically observed variation of the ratio between the keeper-orifice region pressure and the total, upstream pressure. Good agreement is obtained with our control-volume based approach and numerical results. The statistical analysis we conducted on the numerical data demonstrated that (i) the pressure within the keeper-orifice region is most likely 4% of the total, upstream pressure, although cases for which the flow remains subsonic within the keeper-orifice region result in much higher pressure recovery values (up to 80%), and (ii) $1/\beta$, the ratio between the cathode pressure-diameter product, $P_{T}d_{c}$, and the keeper-orifice region pressure-distance ratio, $P_{ko}D_{ko}$, is most likely $\approx 40$ and is related to the ratio of ignition to nominal mass flow rates.
@@
Our study is limited to that of the neutral, noble gas, flow in cathodes and does not consider any plasma effects (e.g., neutral heating due to charge-exchange collisions) that may alter the flow structure. However, the neutral flow field obtained from DSMC may be used as a first guess in a coupled plasma / neutral numerical model (e.g., PIC/DSMC). The study is further limited by the data upon which it is based: we considered a wide, albeit non-exhaustive, range of rescaled parameters describing orificed cathodes with an enclosed keeper, and our conclusions may not apply to cathode configurations that are not covered by our dataset.
@@
The study demonstrates that the flow within the orifice, prior to cathode ignition, is more accurately represented by a Fanno flow approach than an isentropic flow or Poiseuille flow one. The Poiseuille flow model should not be used to estimate the pressure (and/or density and temperature) within hollow cathodes because most assumptions upon which the it is based are invalid in the regime in which cathodes operate. Finally, as shown through our statistical analysis and explained by simple conservation laws, a fifty-fold increase in mass flow rate is likely to result in a minimum voltage in the Paschen-like, DC discharge. Although we only considered xenon, these conclusions are likely applicable to other noble, monatomic gases such as argon.
@@ Acknowledgments
This work was conducted at the Princeton Collaborative Research Facility supported by the U.S. Department of Energy through contract DE-AC02-09CH11466. The computational resources of the United States Naval Academy are gratefully acknowledged.
@@ V.2.1 Continuity
$\dfrac{P_{o}}{P_{k}}=\dfrac{M_{k}}{M_{o}}\sqrt{\dfrac{T_{o}}{T_{k}}}\bar{d}^{2},$ (20)
@@ V.2.2 Energy
$\dfrac{T_{o}}{T_{k}}=\dfrac{1+\dfrac{\gamma-1}{2}M_{k}^{2}}{1+\dfrac{\gamma-1}{2}M_{o}^{2}}.$ (21)
@@ V.2.3 Momentum along $z$-axis
$\begin{split}-\left[\left(\rho_{o}u_{o}^{2}+P_{o}\right)\pi r_{o}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{o}^{2}\right)\right]\\ +\left[\left(\rho_{k}u_{k}^{2}+P_{k}\right)\pi r_{k}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{k}^{2}\right)\right]&\\ +\Pi_{rz}2\pi D_{ko}r_{e}&=0.\end{split}$ (22)
@@
$\begin{split}\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{k}}{P_{o}}\cdot\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{k}^{2}\right)&\\ -\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{o}^{2}\right)&+\dfrac{\left<\tau\right>2D_{ko}r_{e}}{P_{T}ro^{2}}.\end{split}$ (23)
Inserting Equations 20 and 21 into Equation 23, and, because $r_{e}\approx 2r_{k}$, we have
$\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{o}}{P_{T}}\mathcal{F}\left(M_{o},M_{k}\right)+8\dfrac{\left<\tau\right>}{P_{T}}\bar{d}\dfrac{D_{ko}}{d_{o}},$ (24)

### v2

@@ IV.1.2 Subsonic flow
In some cases where $d_{o}=d_{k}$, the keeper induces a (static) pressure rise that is large enough to prevent the orifice from choking. In this configuration the flow remains subsonic throughout the orifice-keeper region, and the supersonic expansion only occurs in the keeper orifice (as opposed to the main cathode orifice). Because the flow is subsonic in this particular case, the keeper-to-total pressure ratio, $P_{ko}/P_{T}$, is higher than that in the supersonic case. Figures 22 and 23 show a contour of the Mach number and static pressure for the subsonic case, respectively. For high-enough mass flow rates, however, the flow may be supersonic. The exact point at which the flow transitions from subsonic to supersonic for the case where $d_{o}=d_{k}$ is beyond the scope of this article.
Some heaterless cathodes may have a configuration for which $d_{k}<d_{o}$ (i.e., $d_{k}/d_{o}<1$). This configuration is beyond the scope of this article: a continuum or near-continuum solver is likely more appropriate for this case as the flow will remain subsonic and in the continuum regime throughout the orifice-keeper region. This case is illustrated in Figure 24, which shows the total pressure and Mach number for the case where $d_{k}/d_{o}=0.5$, with $d_{o}/d_{c}=0.3$ and $D_{ko}/d_{o}=3$. The fluid solver (ANSYS-CFX) was used for this particular case. The total pressure remains approximately constant ($\approx 1$ Torr) in the cathode channel and the orifice-keeper region. Much like the case $d_{o}=d_{k}$, the supersonic expansion occurs in the keeper orifice.
Figure 18: Centerline values for $d_{o}/d_{c}=0.1$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$. A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case. The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
Figure 19: Centerline values for $d_{o}/d_{c}=0.3$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$. A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case. The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
Figure 20: Centerline values for $d_{o}/d_{c}=0.5$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$. A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case. The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
Figure 21: Centerline values for $d_{o}/d_{c}=0.7$ and $P_{T}d_{c}=1$ Torr-cm. (a,d,g) Normalized density, (b,e,h) normalized pressure, (c,f,i) Mach number. Density and pressure are normalized by the upstream values ($n_{T}=6.4\cdot 10^{22}$ m-3 and $P_{T}=267$ Pa, respectively). The horizontal axis is normalized by the orifice-keeper distance, $D_{ko}$. A scaled schematic of each configuration investigated is shown under each $d_{k}/d_{o}$ case. The keeper entrance for $D_{ko}/d_{o}=1$ overlaps the keeper exit for $D_{ko}/d_{o}=0.5$
@@
Figure 24: (a) Total pressure and (b) Mach number for the case $d_{o}/d_{c}=0.3$, $D_{ko}/d_{o}=3$, and $d_{k}/d_{o}=0.5$.
@@ V Conclusion
We have performed a parametric study to quantify the effect of the changes in keeper geometrical parameters on the neutral flow field within hollow cathodes. Because cathode ignition occurs typically between the keeper and orifice plates, we focused our study on the keeper-orifice region. Using the DSMC method, we showed that the neutral cathode flow results in most cases in an underexpanded, supersonic, rarefied jet with Knudsen numbers above 0.1, along with a toroidal subsonic recirculation zone. The recirculation zone is bounded by the underexpanded jet viscous layer, the cathode orifice plate, and the keeper plate. We have shown that simple conservation laws can explain the experimentally and numerically observed variation of the ratio between the keeper-orifice region pressure and the total, upstream pressure. Good agreement is obtained with our control-volume based approach and numerical results. The statistical analysis we conducted on the numerical data demonstrated that (i) the pressure within the keeper-orifice region is most likely 4% of the total, upstream pressure, although cases for which the flow remains subsonic within the keeper-orifice region result in much higher pressure recovery values (up to 80%), and (ii) $1/\beta$, the ratio between the cathode pressure-diameter product, $P_{T}d_{c}$, and the keeper-orifice region pressure-distance ratio, $P_{ko}D_{ko}$, is most likely $\approx 50$ (excluding the case where $d_{k}=d_{o}$) and is related to the ratio of ignition to nominal mass flow rates.
@@
Our study is limited to that of the neutral xenon flow in cathodes and does not consider any plasma effects (e.g., neutral heating due to charge-exchange collisions) that may alter the flow structure. However, the neutral flow field obtained from DSMC may be used as a first guess in a coupled plasma / neutral numerical model (e.g., PIC/DSMC). The study is further limited by the data upon which it is based: we considered a wide, albeit non-exhaustive, range of rescaled parameters describing orificed cathodes with an enclosed keeper, and our conclusions may not apply to cathode configurations that are not covered by our dataset.
@@
The study demonstrates that the flow within the orifice, prior to cathode ignition, is more accurately represented by a Fanno flow approach than an isentropic flow or Poiseuille flow one. The Poiseuille flow model should not be used to estimate the pressure (and/or density and temperature) within hollow cathodes because most assumptions upon which the it is based are invalid in the regime in which cathodes operate. Select numerical cases also show that fluid solvers, while offering qualitative agreement with DSMC results in the keeper-orifice region, differ quantitatively from DSMC in regions where the gas becomes transitional and rarefied. Finally, as shown through our statistical analysis and explained by simple conservation laws, a fifty-fold increase in mass flow rate is likely to result in a minimum voltage in the Paschen-like, DC discharge. Although we only considered xenon, these conclusions are likely applicable to other noble, monatomic gases such as argon: while a lighter gas will have a lower mass flow rate for the same total pressure, temperature, and geometry (see Equation 11), all other scaled parameters (Mach number, Knudsen number, pressure-diameter product, ratio of specific heats, geometric parameters) remain similar.
@@ Acknowledgments
The views expressed in this article are those of the authors and do not reflect the official policy or position of the U.S. Naval Academy, Department of the Navy, the Department of Defense, the U.S. Government, or any agency thereof. This work was conducted at the Princeton Collaborative Research Facility supported by the U.S. Department of Energy through contract DE-AC02-09CH11466. The computational resources of the United States Naval Academy are gratefully acknowledged.
@@ V.2.1 Continuity
$\dfrac{P_{o}}{P_{k}}=\dfrac{M_{k}}{M_{o}}\sqrt{\dfrac{T_{o}}{T_{k}}}\bar{d}^{2},$ (B1)
@@ V.2.2 Energy
$\dfrac{T_{o}}{T_{k}}=\dfrac{1+\dfrac{\gamma-1}{2}M_{k}^{2}}{1+\dfrac{\gamma-1}{2}M_{o}^{2}}.$ (B2)
@@ V.2.3 Momentum along $z$-axis
$\begin{split}-\left[\left(\rho_{o}u_{o}^{2}+P_{o}\right)\pi r_{o}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{o}^{2}\right)\right]\\ +\left[\left(\rho_{k}u_{k}^{2}+P_{k}\right)\pi r_{k}^{2}+P_{ko}\pi\left(r_{e}^{2}-r_{k}^{2}\right)\right]&\\ +\Pi_{rz}2\pi D_{ko}r_{e}&=0.\end{split}$ (B3)
@@
$\begin{split}\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{k}}{P_{o}}\cdot\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{k}^{2}\right)&\\ -\dfrac{P_{o}}{P_{T}}\bar{d}^{2}\left(1+\gamma M_{o}^{2}\right)&+\dfrac{\left<\tau\right>2D_{ko}r_{e}}{P_{T}ro^{2}}.\end{split}$ (B4)
Inserting Equations B1 and B2 into Equation B4, and, because $r_{e}\approx 2r_{k}$, we have
$\dfrac{P_{ko}}{P_{T}}\left(\bar{d}^{2}-1\right)=\dfrac{P_{o}}{P_{T}}\mathcal{F}\left(M_{o},M_{k}\right)+8\dfrac{\left<\tau\right>}{P_{T}}\bar{d}\dfrac{D_{ko}}{d_{o}},$ (B5)